OpenStudy (anonymous):

Find the equation of the circle whose center and radius are given. center ( 7, -3), radius = 7

OpenStudy (anonymous):

Equation of a circle is given in the form \[(x-h)^2+(y-k)^2=r^2\] where (h,k) is your center and r is the radius. If you would like a derivation of the formula, please ask :)

OpenStudy (anonymous):

So the equation would be (x-7)^2+(y- -3)^2=7^2?